OpenStudy (anonymous):

help please:( Find the Taylor series for f(x) centered at the given value of a. [Assume that f has a power series expansion. Do not show that R(x) → 0.] for 8cos(x) at a=3pi

OpenStudy (dan815):

hi

OpenStudy (anonymous):

hello!

OpenStudy (dan815):

taylor series is off linear-->quadratic-->cub---->quartic and so on approximations to hte next point

OpenStudy (anonymous):

right, i got that. however once i expand the taylor series my problem is simplifying it to a summation

OpenStudy (dan815):

okay show me what u got

OpenStudy (dan815):

http://prntscr.com/57u0nc

OpenStudy (camerondoherty):

Dan will teach you He's the best

OpenStudy (dan815):

okay so basically u can decompose the sin and cos

OpenStudy (dan815):

and the coeffeicients infront of them as the summation of odd a even increments of n factorial

OpenStudy (anonymous):

so far i've gotten, \[\frac{ -8(x-3\pi)^0} { 0! }+\frac{ 8(x-3\pi)^1} { 1! }-\frac{ 8(x-3\pi)^2} { 2! }+\frac{8(x-3\pi)^3} { 3! }\] and so on.. what do i do from here?

OpenStudy (camerondoherty):

He's #2kewl4skewl

OpenStudy (dan815):

okay eyah thats good

OpenStudy (dan815):

so now rewrite that as a summation alterating negative sign so

OpenStudy (dan815):